# SAFREE: Training-Free and Adaptive Guard for Safe Text-to-Image And Video Generation

### v1

@@ A.2 Baselines and Evaluation Metrics
Baselines. We primarily compare our method with recently proposed training-free approaches allowing instant weight editing or filtering: variants of SLD (Schramowski et al., 2023) and UCE (Gandikota et al., 2024) and RECE (Gong et al., 2024). UCE and RECE are training-free weight modification methods with a closed-form solution. We note that the editing time of RECE largely varies according to the concept to remove (e.g., several seconds to minutes). In addition, we compare SAFREE with training-based baselines to highlight the advantages of our approach encompassing decent safeguard capability through a training-free framework: ESD (Gandikota et al., 2023), SA (Heng & Soh, 2023), CA (Kumari et al., 2023), MACE (Lu et al., 2024), SDID (Li et al., 2024b).
Hyperparameters. Following Si et al. (2024), we use two different values for $s$: $s_{1}=0.9$ and $s_{2}=0.2$ for the first and second stage blocks of the decoder, respectively. $\alpha=0.01$, $\gamma=10$ throughout the paper.
@@
Evaluation Metrics. We measure the Attack Success Rate (ASR) on adversarial prompts in terms of nudity following Gong et al. (2024) to evaluate the safeguard capability of methods. To evaluate the original generation quality of safe generation or unlearning methods, we measure the FID (Heusel et al., 2017), CLIP score, and a fine-grained faithfulness evaluation metric TIFA score (Hu et al., 2023) on COCO-30k (Lin et al., 2014) dataset. Among these, we randomly select 1k samples for evaluating FID and TIFA. In artist concept removal tasks, we use LPIPS (Zhang et al., 2018) to calculate the perceptual difference between SD-v1.4 output and filtered images following Gong et al. (2024). To more accurately evaluate whether the model removes characteristic (artist) ”styles” in its output while preserving neighbor and interconnected concepts, we frame the task as a Multiple Choice Question Answering (MCQA) problem. Given the generated images, we ask GPT-4o (gpt 4o, 2024) to identify the best matching artist name from five candidates. For safe Text-to-Video evaluation metrics, we follow the automatic evaluation via ChatGPT proposed by T2VSafetybench (Miao et al., 2024). We input sampled 16 video frames along with the same prompt design presented in T2VSafetybench to GPT-4o (gpt 4o, 2024) for binary safety checking.
Appendix B More Qualitative Visualization
@@
We provide more visualization in this Appendix. We provide visualization of artist concept removal in Figs. 5, 6, 7 and 8, where we remove ’Van Gogh’ in the model. Across  Figs. 5, 6 and 7, we observe that our SAFREE can effectively remove ’Van Gogh’ without updating any model weights while other methods, even for training-based method, still struggle for removing this concept. Meanwhile, SAFREE preserves the quality of the desirable concepts in the given prompts. SAFREE can generate the same subjects/scenes as the base model did but remove the targeted style concepts. Furthermore, as shown in Fig. 8, we test both SAFREE and other baseline methods with text prompts containing other artist concepts. All models removed the ’Van Gogh’ concept in their own way. SAFREE successfully preserved other artist styles by maintaining a high similarity to the original SD-1.4 outputs. Meanwhile, other methods like CA/SLD failed to hold the desirable concept. We further show more results by removing the ’nudity’ concept in Figs. 9, 10 and 11, and draw a similar conclusion.
@@
We further change our diffusion model backbones to more advanced SDXL (Podell et al., 2023) and SD-v3 (stabilityai, 2024), as well as Text-to-Video generation backbone models, ZeroScopeT2V (zeroscope, 2024) and CogVideoX (Yang et al., 2024d). As shown in Fig. 12, our method shows robustness across Text-to-Image model backbones, and can effectively filter user-defined unsafe concepts but still preserve the safe concepts in the given toxic prompts. As illusrated in Figs. 13, 14, 15, 17, 18 and 19, our SAFREE shows good generalization ability to Text-to-Video settings. It helps to guard against diverse unsafe/toxic concepts (e.g., animal abuse, porn, violence, terrorism) while preserving quality to the remaining desirable content (e.g., building/human/animals).
Appendix C Limitation & Broader Impact
@@
While our SAFREE demonstrates remarkable effectiveness in concept safeguarding and generalization abilities across backbone models and tasks. We notice that it is still not a perfect method to ensure safe generation in any case. Specifically, our filtering-based SAFREE method exhibits limitations when toxic prompts become much more implicit and in a chain-of-thought style or under an Image-to-Video settting (Yu et al., 2024). such kind of toxic prompts can still jailbreak our SAFREE and yield unsafe/inappropriate content generation. However, we also note that perfect safeguarding in generative models is a challenging open problem that needs more future studies.
@@
Appendix D License Information
@@
Figure 12: More T2I generated examples with different diffusion model backbone (SDXL and SD-v3). Our SAFREEcan guard the ’nudity’ concept in any given diffusion models and still preserve safe concepts in the toxic prompts. We manually masked unsafe generated results for display purposes.

### v2

@@ A.2 Baselines and Evaluation Metrics
Baselines. We primarily compare our method with recently proposed training-free approaches allowing instant weight editing or filtering: variants of SLD (Schramowski et al., 2023) and UCE (Gandikota et al., 2024) and RECE (Gong et al., 2024). In addition, we compare SAFREE with training-based baselines to highlight the advantages of our approach encompassing decent safeguard capability through a training-free framework: ESD (Gandikota et al., 2023), SA (Heng & Soh, 2023), CA (Kumari et al., 2023), MACE (Lu et al., 2024), SDID (Li et al., 2024b). We provide further details of baselines in the Appendix.
@@
Evaluation Metrics. We measure the Attack Success Rate (ASR) on adversarial prompts in terms of nudity following Gong et al. (2024) to evaluate the safeguard capability of methods. To evaluate the original generation quality of safe generation or unlearning methods, we measure the FID (Heusel et al., 2017), CLIP score, and a fine-grained faithfulness evaluation metric TIFA score (Hu et al., 2023) on COCO-30k (Lin et al., 2014) dataset. Among these, we randomly select 1k samples for evaluating FID and TIFA. In artist concept removal tasks, we use LPIPS (Zhang et al., 2018) to calculate the perceptual difference between SD-v1.4 output and filtered images following Gong et al. (2024). To more accurately evaluate whether the model removes characteristic (artist) ”styles” in its output while preserving neighbor and interconnected concepts, we frame the task as a Multiple Choice Question Answering (MCQA) problem. Given the generated images, we ask GPT-4o (gpt 4o, 2024) to identify the best matching artist name from five candidates.
Appendix B Analysis: Correlation Between Undesirable Prompts and Distance to Toxic Concept Subspace.
Refer to caption Refer to caption
Appendix C More Qualitative Visualization
@@
We provide more visualization in this Appendix. We provide visualization of artist concept removal in Figs. 5, 6, 7 and 8, where we remove ’Van Gogh’ in the model. Across  Figs. 5, 6 and 7, we observe that SAFREE can effectively remove ’Van Gogh’ without updating any model weights while other methods, even for training-based method, still struggle for removing this concept. Meanwhile, SAFREE keep maximum faithfulness to the desirable concepts in the given prompts. SAFREE can generate the same subjects/scenes as the base model did but remove the targeted style concepts. Furthermore, as shown in Fig. 8, we test both SAFREE and other baseline methods with text prompts containing other artist concepts. All models removed the ’Van Gogh’ concept in their own way. Ours successfully preserved other artist styles by maintaining a high similarity to the original SD-1.4 outputs. Meanwhile, other methods like CA/SLD failed to hold the desirable concept. We further show more results by removing the ’nudity’ concept in Figs. 9, 10 and 11, and draw a similar conclusion.
@@
We further change our diffusion model backbones to more advanced SDXL (Podell et al., 2023) and SD-v3 (stabilityai, 2024), as well as Text-to-Video generation backbone models, ZeroScopeT2V (zeroscope, 2024) and CogVideoX (Yang et al., 2024c). As shown in Fig. 12, our method shows robustness across Text-to-Image model backbones, and can effectively filter user-defined unsafe concepts but still keep maximum faithfulness to the safe concepts in the given toxic prompts. As illusrated in Figs. 13, 14, 15, 17, 18 and 19, SAFREE shows good generalization ability to Text-to-Video settings. It helps to guard against diverse unsafe/toxic concepts (e.g., animal abuse, porn, violence, terrorism) while preserving faithfulness to the remaining desirable content (e.g., building/human/animals).
Appendix D Extra Discussion with Image Attribute Control Works
Given that our method perturbs the text embedding fed into the model, we further discuss related works that utilize text embedding modifications for enhancing performance.
Baumann et al. (2024) propose optimization-based and optimization-free methods in the CLIP text embedding space for fine-grained image attribute editing (e.g., age). Unlike their focus on image attribute editing, SAFREE manipulates text embeddings for safe text-to-image/video generation.
Zarei et al. (2024) improve attribute composition in T2I models via text embedding optimization. In contrast, SAFREE is training-free and not only perturbs text embeddings but also re-attends latent space in T2I/T2V models for safe generation.
Appendix E Limitation & Broader Impact
@@
While SAFREE demonstrates remarkable effectiveness in concept safeguarding and generalization abilities across backbone models and tasks, we notice that it is still not a perfect method to ensure safe generation in any case. Specifically, our filtering-based SAFREE method exhibits limitations when toxic prompts become much more implicit and in a chain-of-thought style. such kind of toxic prompts can still jailbreak SAFREE and yield unsafe/inappropriate content generation. However, we also note that perfect safeguarding in generative models is a challenging open problem that needs more future studies.
@@
Appendix F License Information
@@
Figure 12: More T2I generated examples with different diffusion model backbone (SDXL and SD-v3). SAFREEcan guard the ’nudity’ concept in any given diffusion models and still keep faithfulness to the safe concepts in the toxic prompts. We manually masked unsafe generated results for display purposes.